2. Residents in City X tend to read Newspaper A or Newspaper B, or not read any newspaper at all. Statistics shows that, after a year . For people reading newspaper A, 30% of them will switch to reading newspaper B and For people reading newspaper B, 20% of them will switch to reading newspaper A and . For people not reading newspaper, 20% will switch to reading newspaper A and 10% to 10% to not reading newspaper 10% to not reading newspaper. reading newspaper B. The current percentage of residents reading Newspaper A, Newspaper B, and not reading any newspaper are 40%, 50% and 10%, respectively. By hand calculation, what are the percentages of readership after 1 year? Then calculate the expected percentages of readership after 5, 10, 15 and 20 years (you may use any computational software). Do the percentages tend to steady state values? matrix
Answer #1

Let total number of citizens in City X in 0th year is N.

Let number of citizens who read newspaper A in 0th year is Aj.

Let number of citizens who read newspaper B in 0th year is Bj.

Let number of citizens who do not read any newspaper in 0th year is Cj.

j=0,1,2,....

Given:- A0=0.4N , B0=0.5N , C0=0.1N Rightarrow A0/N=0.4 , B0/N=0.5 , C0/N=0.1

According to the problem :-

A1 = 0.6A0 + 0.2B0 + 0.2C0

B1 = 0.3A0 + 0.7B0 + 0.1C0

C1 = 0.1A0 + 0.1B0 + 0.7C0

herefore  A1/N = (0.6A0 + 0.2B0 + 0.2C0)/N = 0.6A0/N + 0.2B0/N + 0.2C0/N = 0.6imes0.4 + 0.2imes0.5 + 0.2imes0.1 = 0.36 = 36%

herefore B1/N = (0.3A0 + 0.7B0 + 0.1C0)/N = 0.3A0/N + 0.7B0/N + 0.1C0/N = 0.3imes0.4 + 0.7imes0.5 + 0.1imes0.1 = 0.48 = 48%

herefore C1/N = (0.1A0 + 0.1B0 + 0.7C0)/N = 0.1A0/N + 0.1B0/N + 0.7C0/N = 0.1imes0.4 + 0.1imes0.5 + 0.7imes0.1 = 0.16 = 16%

The following calculations are done using the following R code :-

> N-100 A 0 > for (i in 1:20) > #A [1] denotes number of citizens who read newspaper A in Oth year > #B [1] denotes number of citizens who read newspaper B in 0th year > #C[1] denotes number of citizens who do not read any newspaper in Oth year > #A [1+1] denotes number of citizens who read newspaper A in i-th year > #B [1+1] denotes number of citizens who read newspaper B in i-th year > #Ci+1] denotes number of citizens who do not read any newspaper in 1-th year > A [1] 40.00000 36.00000 34.40000 33.76000 33.50400 33.40160 33.36064 33.34426 [91 33.33770 33.33508 33.33403 33.33361 33.33345 33.33338 33.33335 33.33334 [17] 33.33334 33.33333 33.33333 33.33333 33.33333 [1] 50.00000 48.00000 46.00000 44.48000 43.44000 42.76480 42.33920 42.07565 [9] 41.91424 41.81608 41.75667 41.72081 41.69921 41.68621 41.67840 41.67371 [17] 41.67090 41.66920 41.66819 41.66758 41.66722 [1] 10.00000 16.00000 19.60000 21.76000 23.05600 23.83360 24.30016 24.58010 [9] 24.74806 24.84883 24.90930 24.94558 24.96735 24.98041 24.98825 24.99295 [17] 24.99577 24.99746 24.99848 24.99909 24.99945

For N=100, we see A[6]=33.40160, B[6]=42.76480, C[6]=23.83360

herefore After 5 years A5/N=33.40160/100=33.40106% , B5/N=42.76480/100=42.7948% , C5/N=23.83360/100=23.8336%

For N=100, we see A[11]=33.33403, B[11]=41.75667, C[11]=24.90930

herefore After 10 years A10/N=33.33403/100=33.33403% , B10/N=41.75667/100=41.75667% , C10/N=24.90930/100=23.8336%

For N=100, we see A[16]=33.33334, B[16]=41.67371, C[16]=24.99295

herefore After 15 years A15/N=33.33334/100=33.33334% , B15/N=41.67371/100=41.67371% , C15/N=24.99295/100=24.99295%

For N=100, we see A[21]=33.33333, B[21]=41.66722, C[21]=24.99945

herefore After 20 years A20/N=33.33333/100=33.33333% , B20/N=41.66722/100=41.66722% , C20/N=24.99945/100=24.99945%